how many kilograms of pure salt must be added to 20 kilograms of a 10 percent salt solution to obtain a 25 percent salt solution?
Answer by checkley71(8403)

20*.1+X=.25(20+X) OR 2+X=5+.25X OR .75X=3 OR X=3/.75 OR 4 KILOGRAMS OF PUE SALT NEEDS TO BE ADDED
PROOF 20*.1+4=.25(20+4) OR 2+4=5+1 OR 6=6